4. 4.1% at 630 nm solution . What is the absorbance of the solution? of blue dye (-1.38x 105 cm rmofi L) has a %T of 3 7. B. What i the concentration of the blue dye? olve the equation, p x(b) for x and y (via substitution) given: A, p = 6.84, b = 3.7 B. p 1.71, b 1.85?
Answer #1

2)

A)

Answer

0.7053M

Explanation

Molarity = no of moles per liter of solution

No of moles of NaOCl = mass of NaOCl/molar mass of NaOCl

No of moles of NaOCl = 5.25g/74.44g/mol

No of moles of NaOCl = 0.07053

Therefore,

molarity of NaOCl = (0.07053mol/100ml)×1000ml = 0.7053M

B)

Answer

0.0973M

Explanation

Dilution formula is

C1 × V1 = C2 × V2

C1 = initial concentration

V1 = Initial Volume

C2 = Final concentration

V2 = Final Volume

C2 = C1×V1/V2

C2 = 4ml × 0.7053M/29ml

C2 = 0.0973M

3)

Answer

2.0609g

Explanation

Molarity of solution 2A = 0.7053M

Molar mass of NaCl = 58.44g/mol

mass of NaCl required for 1L of solution= 58.44g/mol × 0.7053mol = 41.218g

mass of NaCl required for 50ml solution = (41.218g/1000ml)×50ml = 2.0609g